OpenStudy (anonymous):

Medal For Help _ Find The Equation Of The LIne With A slope of - 4/3 and a y- intercept of (0,1) a. 4x +3y =3 b.-4x + 3y=3 c. 3x+4y=4 d.3x-4y=4 I'm getting c but not 100% sure

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Ok. Show me your work. So I can tell you what you did wrong.

OpenStudy (anonymous):

Find The Equation Of The Line With A slope of - 4/3 and a y- intercept of (0,1) here we need to use the slope point form of the line (y-y1)= m(x-x1) here we have the first point is ( x1,y1) = (0,1) ans slope = -4/3 so equation is ( y - 1) =(-4/3 ) ( x-0 ) simplifying we get y-1 = (-4/3) x 3y-3 = -4x 4x + 3y = 3 this is the equation of the line...

OpenStudy (anonymous):

So it a ?

OpenStudy (anonymous):

If it's a I was way off

OpenStudy (anonymous):

yes this this is" a" Got it now ?

OpenStudy (anonymous):

Yeap I got it when u broke it down.I see where I made a error. TY